OpenStudy (cm1070):

Waylan has two cardboard boxes that are rectangular prisms. Each cardboard box has the same height, h, but the square bases of each cardboard box have different side measurements, w. He wants to find the amount of space, S, inside each box. Considering that each box has a square base, he decides to use the formula below for his calculation. The square base on Box A has a side measurement of 2 feet, and the square base on Box B has a side measurement of 0.75 of a foot. Each box has a height of 4 feet. Box A contains cubic feet of space, and Box B contains cubic feet of space. Box

OpenStudy (phi):

the volume of a prism is area of base * height or using letters: V= w*w * h where the base is a square whose sides are equal to w

OpenStudy (phi):

you are told The square base on Box A has a side measurement of 2 feet and Each box has a height of 4 feet. the volume of Box A is 2*2*4= 16 the volume of Box B is ¾ * ¾ * 4 = 9/4 = 2.25